A pilot flew a jet from Montreal to Los Angeles, a distance of 2000mi. on the return trip the average speed was 20% faster than the but-bound speed. The round-trip look 9h to 10 min. what was the speed from Montreal to Los Angeles

Answer by mananth(16949)

A pilot flew a jet from Montreal to Los Angeles, a distance of 2000mi. on the return trip the average speed was 20% faster than the but-bound speed. The round-trip look 9h to 10 min. what was the speed from Montreal to Los Angeles
...
forward trip speed = x mph
return trip = 1.2x
distance is the same=2000
...
time forward + time return = 9 1/6 hours= 55/6 hours
...
2000/x + 2000/1.2x = 55/6
1.2*2000 + 2000 /1.2x= 55/6
2400+2000 /1.2x= 55/6
4400/1.2x= 55/6
x= 6*4400 /55*1.2
distance = 400 miles
...
m.ananth@hotmail.ca